OpenStudy (anonymous):

Anybody on that can help me with an algebra 2 question??

OpenStudy (amistre64):

post it and we shall see

OpenStudy (anonymous):

Okay! :)

OpenStudy (amistre64):

seems like a big problem, is it gonna fit on here?

OpenStudy (anonymous):

The pH of a liquid is a measure of how acidic or basic it is. The concentration of hydrogen ions in a liquid isalabeled, h+, use the formula ph=-log{h+} to find the ph level, to the nearest tenth, of a liquid with H+ about 6.2 x 10^-14

OpenStudy (anonymous):

-14.8 14.8 13.2 -13.2

OpenStudy (anonymous):

Its just so late and i am so confused with this math :/

OpenStudy (amistre64):

isnt that just a calculator work? ph = -log(h+), what is ph when h+= 6.2*10^-14

OpenStudy (amistre64):

is your log base 10 or a natural base?

OpenStudy (anonymous):

im not sure cause the ending all it says is 6.2 x 10^-14

OpenStudy (anonymous):

My teacher said all it is, is you plug it in and solve but i do not even know how to plug the stuff where..

OpenStudy (amistre64):

where it says h+, you put in the 6.2 10^(-14)

OpenStudy (amistre64):

log in your case should pertain to base 10 what the means is that log(10) = 1

OpenStudy (anonymous):

so ph= -log (6.2 x 10^-14)

OpenStudy (amistre64):

ph = -log(6.2*10^(-14)) you will need to familiarize yourself with log rules if you have to do this by hand, otherwise just let the calculator do it

OpenStudy (amistre64):

log(ab) = log(a) + log(b) log(k^n) = n log(k)

OpenStudy (anonymous):

I typed that in and it didnt give me a decimal, just a weird fraction

OpenStudy (amistre64):

-log(6.2*10^(-14)) -( log(6.2) + log(10^(-14)) ) -( log(6.2) -14 log(10)) -( log(6.2) -14) -log(6.2) +14 and a calculator gives us: log(6.2) = .7934 so what is 14 - .7924?

OpenStudy (amistre64):

14 - .8 = 13.2 right?

OpenStudy (anonymous):

13.2!!

OpenStudy (amistre64):

then lets go with that one lol

OpenStudy (anonymous):

Do you have time for a few more? Thank you very much!

OpenStudy (amistre64):

this requires you to simply use log rules appropriately, and maybe a calculator at times. but yeah, lets try one more.

OpenStudy (anonymous):

thank you for showing me steps to and not just giving me it straight up,

OpenStudy (anonymous):

Ok, so how would you expand logb 29/61

OpenStudy (anonymous):

i know there is like 4 methods you can use to expand it..

OpenStudy (amistre64):

use paranthesis to clarify the problem, otherwise im not sure what im looking at

OpenStudy (anonymous):

Okie doke one sec

OpenStudy (amistre64):

log_b \(\implies\log_b\)

OpenStudy (anonymous):